Find HCF of 350 and 224 .Also express it as a linear combination of 350 and 224

Grade:10

1 Answers

Arun
5 years ago
Dear Ayushi
 
350 = 1 * 224 + 126
224 = 1 * 126 + 98
126 = 1 * 98 + 28
98 = 3 * 28 + 14
28 = 2 *14 + 0
Hence 14 is HCF.
Now
In the revwrse order try to write 14 as a combination of 350 and 224 only